OpenStudy (anonymous):

What is the solution to the equation 1/ the square root of 8=4^(m-3)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{1}{\sqrt{8}}=4^{(m-3)} }\) like this ?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{8}}= 4^{(m-3)}\]

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{1}{\sqrt{8}}=4^{(m-3)} }\) \(\LARGE\color{blue}{ \frac{1}{\sqrt{8}}=(2^2)^{(m-3)} }\) \(\LARGE\color{blue}{ \frac{1}{\sqrt{8}}=2^{2(m-3)} }\) \(\LARGE\color{blue}{ \sqrt{8}^{~~-1}=2^{2(m-3)} }\) \(\LARGE\color{blue}{ \sqrt{2^3}^{~~-1}=2^{2(m-3)} }\) \(\LARGE\color{blue}{ \sqrt[2]{2^3}^{~~-1}=2^{2(m-3)} }\) \(\LARGE\color{blue}{ (2^{3/2})^{~~-1}=2^{2(m-3)} }\) \(\LARGE\color{blue}{ 2^{~~-1(3/2)}=2^{2(m-3)} }\)

OpenStudy (solomonzelman):

then you get -1(3/2) = 2(m-3)

OpenStudy (solomonzelman):

-3/2 = 2m - 6 6 -3/2 = 2m 12/2 -3/2 = 2m 9/2 = 2m 9/4 = m

OpenStudy (solomonzelman):

Let me know if you need more help.

OpenStudy (anonymous):

ohhh ok I understand it better now thank you.

OpenStudy (solomonzelman):

Anytime :)